ABC Corp Inc. just paid an annual dividend of $2.50 a share. The firm expects future dividends to increase by 3.0% percent annually for the indefinite future. If the required a return on the stock is 10.25%, what is the price of one share of this stock today based on the dividend discount model?
Answer #1

Answer:

D0 = $2.50

G = 3%

Required Return = 10.25%

D1 = D0 * (1 + g)
D1 = $2.50 * (1 + 0.03)
D1 = $2.575

Required Return = D1 / P0 + g
0.1025 = $2.575 / P0 + 0.03
0.1025 – 0.03 = $2.575 / P0
0.0725 = $2.575 / P0
P0 = $2.575 / 0.0725
P0 = $35.52
Answer #2

ANSWER :


Given :


D0 = 2.50 ($)

r = 10.25 % = 0.1025

g = 3 % = 0.03


So,

D1 = D( 1 + 0.34) = 2.50 * 1.03 = 2.575 ($)


Now,


P= D1 / (r - g)  = 2.575 / (0.1025 - 0.03) = 35.52 ($)


So, Price of stock today = $35.52  (ANSWER)



answered by: Tulsiram Garg
Answer #3

P0= DIV1/r-g

= 2.5(1.03)/0.1025-0.03= 35.517

= 35.52

answered by: Ayyas